11. Solve the system by using the substitution method:

y=(x+3)²-1

y=2x+5


A. (-3,-1) and (-1,3)

B. (-1,-3) and (3,-1)

C. (-3,1),(-1,3),(-1,-3) and (3,-1)

D.No solution! it is an inconsistent system


12. Solve the system by using the elimination method:

3x+y²=21

4x²-2y³=-2


A. (2,3) and (2,-3)

B. (-2,3) and (-2,-3)

C. (2,3),(2,-3),(-2,3) and (-2,-3)

D. No solution! It is an Inconsistent System


1
Expert's answer
2022-05-16T10:11:18-0400

11.


"y=2x+5"


"y=2x+5""2x+5=(x+3)^2-1"

Solve the equation


"2x+5=x^2+6x+9-1"

"x^2+4x+3=0"

"(x+3)(x+1)=0"

"x_1=-3, x_2=-1"

Plug solution back into the first equation.


"x_1=-3: y_1=2(-3)+5=-1"

"x_2=-1: y_2=2(-1)+5=3"

A. (-3,-1) and (-1,3)


12.



"3x+y^2=21""4x^3-2y^3=-2"

Given system cannot be solved by using the elimination method.

I think the task is not correct.

If the system is


"3x^2+y^2=21""4x^2-2y^2=-2"

then


"6x^2+2y^2=42""4x^2-2y^2=-2"


"3x^2+y^2=21""4x^2-2y^2+6x^2+2y^2=-2+42"


"3x^2+y^2=21""10x^2=40"


"3(4)+y^2=21""x^2=4"


"y^2=9""x^2=4"

C. (2,3),(2,-3),(-2,3) and (-2,-3)